# Token Hidden Reward: Steering Exploration-Exploitation in Group Relative Deep Reinforcement Learning

### v3

@@ 1 Introduction
The integration of reinforcement learning with verifiable rewards (RLVR) has significantly advanced the reasoning capabilities of large language models (LLMs) (guo2025deepseek; jaech2024openai; team2023gemini). Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) and its variants (i.e., GSPO zheng2025group) have emerged as a widely adopted and empirically successful method for training LLMs on complex reasoning tasks. Models like DeepSeek-R1 (guo2025deepseek), DeepSeek-Math (shao2024deepseekmath), Med-R1 (lai2025med), and Search-R1 (jin2025search) have leveraged GRPO to achieve state-of-the-art performance across diverse domains. Despite these successes, a central and persistent challenge in RL-driven LLM training is managing the inherent exploration-exploitation trade-off (tang2024code; harris2025should). Exploration, sampling uncertain actions to acquire novel information, is crucial for tasks demanding creativity (lu2024llm) and enabling generalization to unseen test cases via scaling algorithms (snell2024scaling). Conversely, exploitation prioritizes optimal decision-making based on current knowledge, a preference in applications requiring high-confidence, low-variance responses, such as medical diagnosis (wu2025medreason). However, effectively shifting the training objective between exploration and exploitation remains an underexplored challenge.
Recent work has begun addressing this pressing challenge through various approaches. chow2024inference examine how to steer the balance between exploration and exploitation via a best-of-$n$ training objective, but their approach relies on an external verifier to select the best candidate among $n$ generations. Contemporaneous works (chen2025pass; mahdavi2025beyond; walder2025pass) introduce Pass@K-training to encourage exploration, though their methods primarily reweight questions based on hardness. Similarly, contemporaneous work (cui2025entropy) steers exploration by controlling entropy, but the analysis is limited to a token’s influence on itself. In parallel, deng2025effect examines the learning dynamics of GRPO, showing how training alters the confidence of correct responses. By downweighting penalties on tokens that reduce this confidence, their method improves greedy decoding performance better exploiting model capabilities. However, their analysis is limited to negative gradients and their role in exploitation.
@@
Motivated by deng2025effect, we examine the intrinsic contribution of each token in the generated responses to the confidence of correct responses and connect this to the exploration–exploitation trade-off. We introduce Token Hidden Reward (THR), a token-level metric that quantifies how individual tokens influence the change in the likelihood of correct responses within the GRPO framework. Our analysis shows that a small subset of tokens carries disproportionately high absolute THR values, while most have negligible impact. Even more interestingly, leveraging the sign of THR, we design a reweighting strategy that explicitly adjusts learning signals : (1) Positive THR tokens amplify the likelihood change of correct responses, strengthening confidence and improving greedy decoding (exploitation); (2) Negative THR tokens preserves probability mass for alternative (than the correct) responses, boosting Pass@K performance (exploration). We specifically compare THR’s token-level reweighting with question-level reweighting approaches such as Pass@K-training, showing that THR provides finer-grained and more effective guidance. Finally, we establish THR’s theoretical and empirical connection to entropy-based exploration methods, while highlighting THR’s efficiency in capturing cross-token interactions. In summary, our main contributions are threefold:
@@ 2 Related Work
Reinforcement Learning for LLM Reasoning. Recent works have explored the use of model-generated solutions as a form of bootstrapping to strengthen the reasoning capabilities of large language models (LLMs)(jaech2024openai; guo2025deepseek; team2025kimi). These methods typically generate candidate solutions using a pre-trained model, then filter them based on intermediate correctness signals(setlur2024rewarding) or final answer correctness (guo2025deepseek; team2025kimi), producing high-quality data to train a new model. Building on the success of reinforcement learning from human feedback (RLHF) (ouyang2022training), follow-up works such as GRPO (shao2024deepseekmath; guo2025deepseek) use online training to further enhance reasoning. Moreover, reinforcement learning directly incorporates the model’s incorrect outputs into training, which has been found to further boost reasoning performance (seed2025seed). Despite these advances, the role of model-generated outputs during training remains underexplored.
Optimizing for inference time objectives. An increasing number of finetuning methods seek to align training with inference-time objectives. Some approaches treat inference-time computation as a flexible post-hoc design choice (snell2024scaling), while others explicitly optimize best-of-$n$ performance during training (huang2025best). The latter, however, depends on an external verifier to select the best output, which complicates scalability. Another direction emphasizes exploitation: deng2025effect reduce penalties on tokens in incorrect responses that positively contribute to correct responses, thereby strengthening the model’s most confident predictions. Their analysis, however, is restricted to negative gradients and does not address exploration. In parallel, several works focus on exploration. Pass@K training (chen2025pass; mahdavi2025beyond; walder2025pass) encourages exploration by reweighting questions based on hardness, but operates only at the question level and overlooks token-level dynamics. Similarly, entropy-based regularization methods such as COV-KL (cui2025entropy) promote exploration by adjusting token entropy, yet they model only a token’s self-influence. By contrast, our work directly targets token-level contributions and cross-token interactions, showing how they govern the exploration–exploitation balance in GRPO.
@@ 3.1 Group Relative Policy Optimization
Group relative policy optimization, introduced in DeepSeek-Math (guo2025deepseek) and DeepSeek-R1 (shao2024deepseekmath), simplifies RLVR by eliminating the value function estimation required in PPO (schulman2017proximal). Instead of learning a separate value network, GRPO computes group-relative rewards within each training batch, reducing training complexity while maintaining stable policy updates. For a query pair $\bm{x}$, the policy $\pi_{\theta}$ samples $G$ responses $\{\bm{y}_{i}\}_{i=1}^{G}$. Each $\bm{y}_{i}$ consists of a sequence of $|\bm{y}_{i}|$ tokens. Given rewards $r_{i}\in\{0,1\}$ for each response, GRPO computes normalized advantages $\hat{A}_{i,k}:=\frac{r_{i}-\mu}{\sigma},$ where $\mu$ and $\sigma$ are the empirical average and standard deviation of the rewards. Specifically for binary rewards $r_{i}\in\{0,1\}$, denoting $q=N^{+}/G$ the fraction of correct ($r_{i}=1$) responses per group, GRPO’s advantage scores become:
@@ 3.2 Likelihood Change of Correct Response in GRPO
A recent study (deng2025effect) analyzed the learning dynamics of GRPO, examining how the likelihood of correct responses $\bm{y}_{i}^{+}$ evolves during training. They proved the following theorem using the unconstrained features framework (yang2017breaking; mixon2022neural; razin2024unintentional):
@@ 4.1 Definition of THR
Exploration is encouraged by a lower increase in the likelihood of correct responses since this preserves some probability mass for alternative outputs.
Exploitation is encourages by a higher increase in the likelihood of correct responses, since this strengthens confidence in those observed correct outputs.
@@ 6 Experiments & Analysis
We evaluate THR’s empirical effectiveness through comprehensive experiments across four dimensions: (1) Demonstrating exploitation ($p>0$) and exploration ($p<0$) capabilities as measured by greedy accuracy and Pass@K performance, (2) comparing our fine-grained token-level control against coarser-grained question-level baselines, (3) analyzing the relationship between THR and prediction entropy, and (4) validating generalizability across a GRPO variant (i.e., GSPO-token (zheng2025group) and Llama architectures.
Experimental settings. For all experiments, we follow zeng2025simplerl and train on the MATH dataset (levels 3–5) (hendrycksmath2021). To accelerate training, we adopt dynamic sampling (yu2025dapo), which discards samples with zero advantage and resamples until a full batch is formed. Unless otherwise specified, all models and methods are fine-tuned with identical reinforcement learning hyperparameters. Specifically, we use four A100 GPUs with a batch size of 256, a learning rate of $1\text{e}^{-6}$, and a mini-batch size of 64, resulting in four updates per step. Training runs for 40 steps, which corresponds to roughly two effective epochs given the higher throughput from dynamic sampling. For each prompt, we generate 8 rollouts. We set the sampling temperature to 1.0, the clipping ratio to 0.2, and the KL loss coefficient to $1\times 10^{-4}$. For the threshold $\tau$, we follow  deng2025effect, defining it as the average influence of the $i^{\prime}$-th correct response’s tokens on the likelihoods of other correct responses. Additional details are provided in Appendix B.
Evaluation setup. Since exploitation focuses on making the best decisions based on existing knowledge (harris2025should), we assess exploitation ability of fine-tuned models by measuring their greedy decoding accuracy. Here we adopt six widely used math benchmarks: three “Hard datasets” (AIME 2025, AIME 2024 (aime_1983_2024), AMC23) and three “Standard datasets” (MATH500 (hendrycksmath2021), Olympiad (he2024olympiadbench), and Minerva Math (lewkowycz2022solving)). To evaluate exploration, we report the unbiased Pass@K@K accuracy (chen2021evaluating) using temperature $1.0$ on the challenging AIME2024, AIME2025 and AMC23 datasets, which require more exploration during attempts. The Pass@K@K metrix is defined as $\text{Pass}@K=\mathbb{E}_{\bm{x}\sim D}\Big[1-{\binom{M-C}{K}}/{\binom{M}{K}}\Big$], where $M\geq K$ is the number of generated responses per question $\bm{x}$, and $C$ denotes the number of correct responses. For all Pass@K@K evaluations, we use $M=256$ and report results for $K=2^{1:8}$.
@@
Qwen2.5-Math-1.5B
Base 0.0 3.3 20.0 7.8 39.6 7.7 24.9 24.1 15.9
@@
Qwen2.5-Math-7B
Base 13.3 6.7 42.5 20.8 64.6 15.8 26.7 35.7 28.3
@@ 6.1 Effectiveness of THR in exploitation and exploration
We use varying-sized Qwen2.5 models (yang2024qwen25mathtechnicalreportmathematical): 0.5B-Ins, Math-1.5B, Math-7B.
@@
Pass@K Training as Question-Level Reweighting. chen2025pass; mahdavi2025beyond; walder2025pass develop RLVR objectives that directly target Pass@K optimization. For GRPO, these amount to re-weightings of the advantage scores in a way that favors “rare successes”—i.e., responses associated with “hard” questions. Crucially, the reweighting is uniform across all tokens and responses for a given question, which we term question-level reweighting. To be concrete, As we show in Appendix D.1, that chen2025pass’s question-level reweighting of vanilla GRPO advantages takes the following simplified form (assuming $G\geq K$):
@@
In practice, we adopt a convex combination $q\cdot\hat{A}_{i,k}+(1-q)\cdot\hat{A}_{i,k}^{@K}$ of vanilla GRPO advantage and the above Pass@K advantage, termed Pass@K-mixed (chen2025pass), to avoid overly suppressing easy questions and preserve valuable learning signals. Empirically, Pass@K-mixed outperforms GRPO on both Qwen2.5-Math-1.5B (Table 3) and Llama3.2-3B-Instruct (Table 9). For training, we use $K=4,G=8$ throughout our experiments.
@@
In summary, both THR and Pass@K training employ what chen2025pass term implicit advantage design to steer exploration. However, THR provides more fine-grained control by operating at the token level, enabling more targeted and effective exploration management.
@@ 6.3 On the Relation of THR with Entropy
In this section, we study the relation between THR and entropy because entropy has long served as a proxy for exploration in RL (wang2018exploration; cui2025entropy).
@@
Dominant tokens overlaps with high entropy tokens. For a confident (low-entropy) token $\mathbf{e}_{\bm{y}_{k^{\prime}}}-\pi(\cdot|\bm{x},\bm{y}_{<k^{\prime}})$ has small magnitude, thus the resulting $\alpha_{\cdot,k^{\prime}}$ in Definition 4.1 tends to be close to zero, leading to a low THR. We analyze the overlap between tokens with high THR scores and those with high entropy. For each sample, we select the same number of high-entropy tokens as high-THR tokens, compute their overlap rate, and plot the kernel density estimate (chen2017tutorial) of the resulting overlap scores in Fig. 3. We find consistently high overlap ratio, often around 90%, indicating a strong correlation between THR and entropy. This finding is consistent with the observation of contemporaneous work (wang2025beyond), demonstrating that training on only the top 20% of high-entropy tokens is sufficient to achieve performance on par with GRPO using all tokens.
@@
Comparison with Cov-KL. cui2025entropy propose Cov-KL as an entropy-based regularization approach focusing on how each token affects the update of itself during training. In contrast, THR, as formalized in Definition 4.1, explicitly captures the cross-token interactions that arise throughout the learning process. As shown in Table 4, THR($p<0$) consistently outperforms Cov-KL in all Pass@K settings, underscoring the importance of modeling cross-token influence for guiding exploration.

### v4

@@ 1 Introduction
The integration of reinforcement learning with verifiable rewards (RLVR) has significantly advanced the reasoning capabilities of large language models (LLMs) (Guo et al., 2025; Jaech et al., 2024; Team et al., 2023). Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its variants (i.e., GSPO Zheng et al. (2025)) have emerged as a widely adopted and empirically successful method for training LLMs on complex reasoning tasks. Models like DeepSeek-R1 (Guo et al., 2025), DeepSeek-Math (Shao et al., 2024), Med-R1 (Lai et al., 2025), and Search-R1 (Jin et al., 2025) have leveraged GRPO to achieve state-of-the-art performance across diverse domains. Despite these successes, a central and persistent challenge in RL-driven LLM training is managing the inherent exploration-exploitation trade-off (Tang et al., 2024; Harris and Slivkins, 2025). Exploration, sampling uncertain actions to acquire novel information, is crucial for tasks demanding creativity (Lu et al., 2024) and enabling generalization to unseen test cases via scaling algorithms (Snell et al., 2024). Conversely, exploitation prioritizes optimal decision-making based on current knowledge, a preference in applications requiring high-confidence, low-variance responses, such as medical diagnosis (Wu et al., 2025). However, effectively shifting the training objective between exploration and exploitation remains an underexplored challenge.
Recent work has begun addressing this pressing challenge through various approaches. Chow et al. (2024) examine how to steer the balance between exploration and exploitation via a best-of-$n$ training objective, but their approach relies on an external verifier to select the best candidate among $n$ generations. Contemporaneous works (Chen et al., 2025; Mahdavi et al., 2025; Walder and Karkhanis, 2025) introduce Pass@K-training to encourage exploration, though their methods primarily reweight questions based on hardness. Similarly, contemporaneous work (Cui et al., 2025) steers exploration by controlling entropy, but the analysis is limited to a token’s influence on itself. In parallel, Deng et al. (2025) examines the learning dynamics of GRPO, showing how training alters the confidence of correct responses. By downweighting penalties on tokens that reduce this confidence, their method improves greedy decoding performance better exploiting model capabilities. However, their analysis is limited to negative gradients and their role in exploitation.
@@
Motivated by Deng et al. (2025), we examine the intrinsic contribution of each token in the generated responses to the confidence of correct responses and connect this to the exploration–exploitation trade-off. We introduce Token Hidden Reward (THR), a token-level metric that quantifies how individual tokens influence the change in the likelihood of correct responses within the GRPO framework. Our analysis shows that a small subset of tokens carries disproportionately high absolute THR values, while most have negligible impact. Even more interestingly, leveraging the sign of THR, we design a reweighting strategy that explicitly adjusts learning signals : (1) Positive THR tokens amplify the likelihood change of correct responses, strengthening confidence and improving greedy decoding (exploitation); (2) Negative THR tokens preserves probability mass for alternative (than the correct) responses, boosting Pass@K performance (exploration). We specifically compare THR’s token-level reweighting with question-level reweighting approaches such as Pass@K-training, showing that THR provides finer-grained and more effective guidance. Finally, we establish THR’s theoretical and empirical connection to entropy-based exploration methods, while highlighting THR’s efficiency in capturing cross-token interactions. In summary, our main contributions are threefold:
@@ 2 Related Work
Reinforcement Learning for LLM Reasoning. Recent works have explored the use of model-generated solutions as a form of bootstrapping to strengthen the reasoning capabilities of large language models (LLMs)(Jaech et al., 2024; Guo et al., 2025; Team et al., 2025). These methods typically generate candidate solutions using a pre-trained model, then filter them based on intermediate correctness signals(Setlur et al., 2024) or final answer correctness (Guo et al., 2025; Team et al., 2025), producing high-quality data to train a new model. Building on the success of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), follow-up works such as GRPO (Shao et al., 2024; Guo et al., 2025) use online training to further enhance reasoning. Moreover, reinforcement learning directly incorporates the model’s incorrect outputs into training, which has been found to further boost reasoning performance (Seed et al., 2025). Despite these advances, the role of model-generated outputs during training remains underexplored.
Optimizing for inference time objectives. An increasing number of finetuning methods seek to align training with inference-time objectives. Some approaches treat inference-time computation as a flexible post-hoc design choice (Snell et al., 2024), while others explicitly optimize best-of-$n$ performance during training (Huang et al., 2025). The latter, however, depends on an external verifier to select the best output, which complicates scalability. Another direction emphasizes exploitation: Deng et al. (2025) reduce penalties on tokens in incorrect responses that positively contribute to correct responses, thereby strengthening the model’s most confident predictions. Their analysis, however, is restricted to negative gradients and does not address exploration. In parallel, several works focus on exploration. Pass@K training (Chen et al., 2025; Mahdavi et al., 2025; Walder and Karkhanis, 2025) encourages exploration by reweighting questions based on hardness, but operates only at the question level and overlooks token-level dynamics. Similarly, entropy-based regularization methods such as COV-KL (Cui et al., 2025) promote exploration by adjusting token entropy, yet they model only a token’s self-influence. By contrast, our work directly targets token-level contributions and cross-token interactions, showing how they govern the exploration–exploitation balance in GRPO.
@@ 3.1 Group Relative Policy Optimization
Group relative policy optimization, introduced in DeepSeek-Math (Guo et al., 2025) and DeepSeek-R1 (Shao et al., 2024), simplifies RLVR by eliminating the value function estimation required in PPO (Schulman et al., 2017). Instead of learning a separate value network, GRPO computes group-relative rewards within each training batch, reducing training complexity while maintaining stable policy updates. For a query pair $\bm{x}$, the policy $\pi_{\theta}$ samples $G$ responses $\{\bm{y}_{i}\}_{i=1}^{G}$. Each $\bm{y}_{i}$ consists of a sequence of $|\bm{y}_{i}|$ tokens. Given rewards $r_{i}\in\{0,1\}$ for each response, GRPO computes normalized advantages $\hat{A}_{i,k}:=\frac{r_{i}-\mu}{\sigma},$ where $\mu$ and $\sigma$ are the empirical average and standard deviation of the rewards. Specifically for binary rewards $r_{i}\in\{0,1\}$, denoting $q=N^{+}/G$ the fraction of correct ($r_{i}=1$) responses per group, GRPO’s advantage scores become:
@@ 3.2 Likelihood Change of Correct Response in GRPO
A recent study (Deng et al., 2025) analyzed the learning dynamics of GRPO, examining how the likelihood of correct responses $\bm{y}_{i}^{+}$ evolves during training. They proved the following theorem using the unconstrained features framework (Yang et al., 2017; Mixon et al., 2022; Razin et al., 2024):
@@ 4.1 Definition of THR
Exploration is encouraged by a lower increase in the likelihood of observed correct responses since this preserves some probability mass for alternative outputs.
Exploitation is encourages by a higher increase in the likelihood of observed correct responses, since this strengthens confidence in those observed correct outputs.
@@ 6 Experiments & Analysis
We evaluate THR’s empirical effectiveness through comprehensive experiments across four dimensions: (1) Demonstrating exploitation ($p>0$) and exploration ($p<0$) capabilities as measured by greedy accuracy and Pass@K performance, (2) comparing our fine-grained token-level control against coarser-grained question-level baselines, (3) analyzing the relationship between THR and prediction entropy, and (4) validating generalizability across a GRPO variant (i.e., GSPO-token (Zheng et al., 2025) and Llama architectures.
Experimental settings. For all experiments, we follow Zeng et al. (2025) and train on the MATH dataset (levels 3–5) (Hendrycks et al., 2021). To accelerate training, we adopt dynamic sampling (Yu et al., 2025), which discards samples with zero advantage and resamples until a full batch is formed. Unless otherwise specified, all models and methods are fine-tuned with identical reinforcement learning hyperparameters. Specifically, we use four A100 GPUs with a prompt batch size of 256 and 8 rollouts per prompt. We use a learning rate of $1\text{e}^{-6}$, and a mini-batch size of 64, resulting in 32 gradient updates per step. Training runs for 40 steps, which corresponds to more than two effective epochs given the higher throughput from dynamic sampling. We set the sampling temperature to 1.0, the clipping ratio to 0.2, and the KL loss coefficient to $1\times 10^{-4}$. For the threshold $\tau$, we follow  Deng et al. (2025, Eq. (8)), defining it as the average influence of the $i^{\prime}$-th correct response’s tokens on the likelihoods of other correct responses. Additional details are provided in Appendix B.
Evaluation setup. Since exploitation focuses on making the best decisions based on existing knowledge (Harris and Slivkins, 2025), we assess exploitation ability of fine-tuned models by measuring their greedy decoding accuracy. Here we adopt six widely used math benchmarks: three “Hard datasets” (AIME 2025, AIME 2024 (Veeraboina, 2023), AMC23) and three “Standard datasets” (MATH500 (Hendrycks et al., 2021), Olympiad (He et al., 2024), and Minerva Math (Lewkowycz et al., 2022)). To evaluate exploration, we report the unbiased Pass@K@K accuracy (Chen et al., 2021) using temperature $1.0$ on the challenging AIME2024, AIME2025 and AMC23 datasets, which require more exploration during attempts. The Pass@K@K metric is defined as $\text{Pass}@K=\mathbb{E}_{\bm{x}\sim D}\Big[1-{\binom{M-C}{K}}/{\binom{M}{K}}\Big$], where $M\geq K$ is the number of generated responses per question $\bm{x}$, and $C$ denotes the number of correct responses. For all Pass@K@K evaluations, we use $M=256$ and report results for $K=2^{1:8}$.
@@
Qwen2.5-Math-1.5B Base 0.0 3.3 20.0 7.8 39.6 7.7 24.9 24.1 15.9
@@
Qwen2.5-Math-7B Base 13.3 6.7 42.5 20.8 64.6 15.8 26.7 35.7 28.3
@@ 6.1 Effectiveness of THR in exploitation and exploration
We use varying-sized Qwen2.5 models (Yang et al., 2024): 0.5B-Ins, Math-1.5B, Math-7B.
@@
Pass@K Training as Question-Level Reweighting. Chen et al. (2025); Mahdavi et al. (2025); Walder and Karkhanis (2025) develop RLVR objectives that directly target Pass@K optimization. For GRPO, these amount to re-weightings of the advantage scores in a way that favors “rare successes”—i.e., responses associated with “hard” questions. Crucially, the reweighting is uniform across all tokens and responses for a given question, which we term question-level reweighting. To be concrete, As we show in Appendix D.1, that Chen et al. (2025)’s question-level reweighting of vanilla GRPO advantages takes the following simplified form (assuming $G\geq K$):
@@
In practice, we adopt a convex combination $q\cdot\hat{A}_{i,k}+(1-q)\cdot\hat{A}_{i,k}^{@K}$ of vanilla GRPO advantage and the above Pass@K advantage, termed Pass@K-mixed (Chen et al., 2025), to avoid overly suppressing easy questions and preserve valuable learning signals. Empirically, Pass@K-mixed outperforms GRPO on both Qwen2.5-Math-1.5B (Table 3) and Llama3.2-3B-Instruct (Table 9). For training, we use $K=4,G=8$ throughout our experiments.
@@
In summary, both THR and Pass@K training employ what Chen et al. (2025) term implicit advantage design to steer exploration. However, THR provides more fine-grained control by operating at the token level, enabling more targeted and effective exploration management.
@@ 6.3 On the Relation of THR with Entropy
In this section, we study the relation between THR and entropy because entropy has long served as a proxy for exploration in RL (Wang et al., 2018; Cui et al., 2025).
@@
Dominant tokens overlaps with high entropy tokens. For a confident (low-entropy) token $\mathbf{e}_{\bm{y}_{k^{\prime}}}-\pi(\cdot|\bm{x},\bm{y}_{<k^{\prime}})$ has small magnitude, thus the resulting $\alpha_{\cdot,k^{\prime}}$ in Definition 4.1 tends to be close to zero, leading to a low THR. We analyze the overlap between tokens with high THR scores and those with high entropy. For each sample, we select the same number of high-entropy tokens as high-THR tokens, compute their overlap rate, and plot the kernel density estimate (Chen, 2017) of the resulting overlap scores in Fig. 3. We find consistently high overlap ratio, often around 90%, indicating a strong correlation between THR and entropy. This finding is consistent with the observation of contemporaneous work (Wang et al., 2025), demonstrating that training on only the top 20% of high-entropy tokens is sufficient to achieve performance on par with GRPO using all tokens.
@@
Comparison with Cov-KL. Cui et al. (2025) propose Cov-KL as an entropy-based regularization approach focusing on how each token affects the update of itself during training. In contrast, THR, as formalized in Definition 4.1, explicitly captures the cross-token interactions that arise throughout the learning process. As shown in Table 4, THR($p<0$) consistently outperforms Cov-KL in all Pass@K settings, underscoring the importance of modeling cross-token influence for guiding exploration.